Miriam is four years older than Ricardo. Ten yeras ago Miriam was 3 times as old as Ricardo was then.
A. How old are they now?
B. Determine whether your answer is correect by checking that it satisfies the conditions of the problem.

Answer by checkley77(12844)

R=X
M=X+4
(X+4)-10=3(X-10)
X+4-10=3X-30
X-6=3X-30
X-3X=-30+6
-2X=-24
X=-24/-2
X=12 AGE OF RICARDO NOW.
12=4=16 AGE OF MIRIAM NOW.
PROOF:
16-10=3(12-10)
6=3*2
6=6